# Soft information decoding with superconducting qubits

### v1

@@ III.2 Decoding data from quantum hardware
While the simulations offer valuable insights, they do not fully capture complexities such as correlated errors, including qubit crosstalk and leakage, which can affect hardware performance. Although recent research has proposed more accurate metrics for closer-to-hardware simulations using calibrated device error rates Hesner et al. (2024), we did not pursue this approach in our current study. Therefore, to better understand the practical implications of soft decoding, we must proceed with experimental implementation on the actual hardware.
@@ III.5 Effect of leakage on the decoding performance
The logical error rate per round shows an apparent increase as the number of rounds grows for hard and soft decoders. The growth rate is less pronounced for the soft decoder than for the hard decoder, but it remains considerable. This trend reveals the substantial effect of leakage on both decoding strategies, even when handling leaked measurement outcomes as maximally ambiguous in soft decoding. Moreover, the observed increase in logical error rate is consistent with the findings from Ref. et al. (2015), who noted a decrease in the $\Lambda$-factor from round to round when implementing repetition codes on hardware.
@@ Appendix A Soft information in other quantum computing platforms
Other quantum computing technologies face similar challenges in achieving accurate qubit state discrimination. For instance, the state measurement in trapped ions involves detecting fluorescence from ions when illuminated with a laser. Ideally, ions in the excited state fluoresce while those in the ground state do not. However, due to background photon count, laser intensity fluctuations, and photodetector efficiency, the resulting signal also forms a distribution of counts. The overlap between distributions for dark state (ground state) and bright state (excited state) similarly leads to a finite misassignment probability Bruzewicz et al. (2019).
In semiconductor quantum dot systems, qubit state readout is often performed through charge sensing, where the tunneling of electrons to and from the quantum dot affects the conductance of a nearby sensor. The conductance signal, influenced by the qubit’s charge state, generates distributions that must be distinguished to infer the qubit state accurately. As with superconducting and ion trap platforms, variations in tunnel rates, sensor sensitivity, and environmental noise contribute to the overlap between these distributions, posing a challenge for reliable qubit state discrimination Burkard et al. (2023).
@@ Appendix D Kernel Density Estimation of $f^{(\bar{z})}(\mu)$
KDE is based on the convolution of a kernel function with the data points. We use the Epanechnikov kernel function as it has a few advantages. Statistically, it is the optimal kernel because it provides the best trade-off between bias and variance. Moreover, it has a compact support, meaning that its value is zero beyond a specific range ($|u|>1$). This property reduces the computational complexity compared to other kernels like the Gaussian kernel since it only accounts for points within its support Epanechnikov (1969). The only free parameter is the kernel bandwidth, which determines the smoothness of the estimated distribution, with a smaller bandwidth leading to a more jagged distribution and a larger bandwidth leading to a smoother distribution. We set the bandwidth for each qubit using cross-validation, splitting the data into a training and a validation set and choosing the bandwidth that minimizes the validation error.
@@ Appendix E Experimental setup
We implement various configurations of the repetition code, encoding both in the $X$ and $Z$ bases and varying the prepared logical state between |$\ket{+z}_{L}$ and |$\ket{-z}_{L}$. We encode the logical state, measure $T$ rounds of stabilizer measurements, and conclude with a final readout of all code qubits in the same basis as initially encoded. We inform our noise model with calibration data, and we utilize Pauli tracing via the Stim software to weigh the edges of the decoding graph Gidney (2021).
We decode the repetition codes using the MWPM algorithm via the PyMatching software Higgott (2021). This algorithm determines the most likely error (MLE) by identifying the minimum-weight perfect matching within the decoding graph. To evaluate the logical error rate, we select the last qubit in the chain—though this selection is arbitrary—to carry the logical information. If the MLE affects this logical qubit, we classify it as a logical error. The results from this process allow us to compare the predicted logical error against the actual logical error observed during the final measurement of the code qubit.
We use a technique known as subsampling et al. (2015); Naveh et al. (2018); Google Quantum AI (2021) to run repetition codes sample-efficiently on hardware. This approach involves deploying the largest possible repetition code that the hardware can accommodate and decoding various subsets to gather data of smaller distances. This strategy generates $n=d-d_{s}+1$ distinct datasets, each corresponding to a repetition code of distance $d_{s}$. An example of subsampling for a repetition code of distance $d_{s}=3$ from a $d=5$ is shown in Fig. 16. By subsampling, we can analyze the performance of repetition codes across multiple distances without the need to individually execute each configuration on the hardware. This method also ensures consistent noise conditions across different experiments, eliminating noise drifts.
Figure 16: Example of subsampling for a repetition code of distance $d_{s}=3$ from a $d=5$ code. The subsampling method allows us to evaluate the performance of repetition codes across multiple distances without the need to individually execute each configuration on the hardware. Figure taken from Google Quantum AI (2021).
@@ Appendix F Simulation setup
To simulate soft decoding of repetition codes as if we run it on the hardware, we first generate outcomes for stabilizer and final code measurements using the circuit-level Pauli noise model. We configure the circuit level noise model with the calibration data obtained from the hardware for two-qubit gates, single-qubit gates, and decoherence times. We use these values alongside the explicit code circuit to estimate the noise model. The Stim software Gidney (2021) facilitates the outcome sampling process by providing tools for Pauli tracing, which we also use to weight the graph.
@@
The software that implements the estimation of outcomes and soft flip probabilities, as well as the software for reweighting the decoding graph, will be made available soon. To ensure rapid execution times, we wrote the performance-critical components in C++. We achieve the kernel density estimation estimation using the MLPack library et al. (2023).

### v2

@@ III.2 Decoding data from quantum hardware
While the simulations offer valuable insights, they do not fully capture complexities such as correlated errors, including qubit crosstalk and leakage, which can affect hardware performance. Although recent research has proposed more accurate metrics for closer-to-hardware simulations using calibrated device error rates [38], we did not pursue this approach in our current study. Therefore, to better understand the practical implications of soft decoding, we must proceed with experimental implementation on the actual hardware.
@@ III.5 Effect of leakage on the decoding performance
The logical error rate per round shows an apparent increase as the number of rounds grows for hard and soft decoders. The growth rate is less pronounced for the soft decoder than for the hard decoder, but it remains considerable. This trend reveals the substantial effect of leakage on both decoding strategies, even when handling leaked measurement outcomes as maximally ambiguous in soft decoding. Moreover, the observed increase in logical error rate is consistent with the findings from Ref. [22], who noted a decrease in the $\Lambda$-factor from round to round when implementing repetition codes on hardware.
@@ Appendix A Soft information in other quantum computing platforms
Other quantum computing technologies face similar challenges in achieving accurate qubit state discrimination. For instance, the state measurement in trapped ions involves detecting fluorescence from ions when illuminated with a laser. Ideally, ions in the excited state fluoresce while those in the ground state do not. However, due to background photon count, laser intensity fluctuations, and photodetector efficiency, the resulting signal also forms a distribution of counts. The overlap between distributions for dark state (ground state) and bright state (excited state) similarly leads to a finite misassignment probability [39].
In semiconductor quantum dot systems, qubit state readout is often performed through charge sensing, where the tunneling of electrons to and from the quantum dot affects the conductance of a nearby sensor. The conductance signal, influenced by the qubit’s charge state, generates distributions that must be distinguished to infer the qubit state accurately. As with superconducting and ion trap platforms, variations in tunnel rates, sensor sensitivity, and environmental noise contribute to the overlap between these distributions, posing a challenge for reliable qubit state discrimination [40].
@@ Appendix D Kernel Density Estimation of $f^{(\bar{z})}(\mu)$
KDE is based on the convolution of a kernel function with the data points. We use the Epanechnikov kernel function as it has a few advantages. Statistically, it is the optimal kernel because it provides the best trade-off between bias and variance. Moreover, it has a compact support, meaning that its value is zero beyond a specific range ($|u|>1$). This property reduces the computational complexity compared to other kernels like the Gaussian kernel since it only accounts for points within its support [41]. The only free parameter is the kernel bandwidth, which determines the smoothness of the estimated distribution, with a smaller bandwidth leading to a more jagged distribution and a larger bandwidth leading to a smoother distribution. We set the bandwidth for each qubit using cross-validation, splitting the data into a training and a validation set and choosing the bandwidth that minimizes the validation error.
@@ Appendix E Experimental setup
We implement various configurations of the repetition code, encoding both in the $X$ and $Z$ bases and varying the prepared logical state between |$\ket{+z}_{L}$ and |$\ket{-z}_{L}$. We encode the logical state, measure $T$ rounds of stabilizer measurements, and conclude with a final readout of all code qubits in the same basis as initially encoded. We inform our noise model with calibration data, and we utilize Pauli tracing via the Stim software to weigh the edges of the decoding graph [25].
We decode the repetition codes using the MWPM algorithm via the PyMatching software [16]. This algorithm determines the most likely error (MLE) by identifying the minimum-weight perfect matching within the decoding graph. To evaluate the logical error rate, we select the last qubit in the chain—though this selection is arbitrary—to carry the logical information. If the MLE affects this logical qubit, we classify it as a logical error. The results from this process allow us to compare the predicted logical error against the actual logical error observed during the final measurement of the code qubit.
We use a technique known as subsampling [22, 42, 20] to run repetition codes sample-efficiently on hardware. This approach involves deploying the largest possible repetition code that the hardware can accommodate and decoding various subsets to gather data of smaller distances. This strategy generates $n=d-d_{s}+1$ distinct datasets, each corresponding to a repetition code of distance $d_{s}$. An example of subsampling for a repetition code of distance $d_{s}=3$ from a $d=5$ is shown in Fig. 16. By subsampling, we can analyze the performance of repetition codes across multiple distances without the need to individually execute each configuration on the hardware. This method also ensures consistent noise conditions across different experiments, eliminating noise drifts.
Figure 16: Example of subsampling for a repetition code of distance $d_{s}=3$ from a $d=5$ code. The subsampling method allows us to evaluate the performance of repetition codes across multiple distances without the need to individually execute each configuration on the hardware. Figure taken from [20].
@@ Appendix F Simulation setup
To simulate soft decoding of repetition codes as if we run it on the hardware, we first generate outcomes for stabilizer and final code measurements using the circuit-level Pauli noise model. We configure the circuit level noise model with the calibration data obtained from the hardware for two-qubit gates, single-qubit gates, and decoherence times. We use these values alongside the explicit code circuit to estimate the noise model. The Stim software [25] facilitates the outcome sampling process by providing tools for Pauli tracing, which we also use to weight the graph.
@@
The software that implements the estimation of outcomes and soft flip probabilities, as well as the software for reweighting the decoding graph, will be made available soon. To ensure rapid execution times, we wrote the performance-critical components in C++. We achieve the kernel density estimation estimation using the MLPack library [43].